Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 10,000 $ 15,000 $ 25,000
Estimated variable manufacturing overhead per machine-hour $ 1.40 $ 2.20
Job P Job Q
Direct materials $ 13,000 $ 8,000
Direct labor cost $ 21,000 $ 7,500
Actual machine-hours used:
Molding 1,700 800
Fabrication 600 900
Total 2,300 1,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

9. What were the company’s predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your answers to 2 decimal places.)

Predetermined Overhead Rate per MH Molding Department Fabrication Department per MH

10. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

moa job p ________

moa job q ________

11. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

moa job p ________

moa job q ________

12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.)

unit product cost ________
Answer #1

9) Molding department = (10000/2500+1.4) = 5.4 per MH

Fabrication overhead = (15000/1500+2.2) = 12.20 per Mh

10) Molding overhead applied

Job P = 1700*5.4 = 9180

Job Q = 800*5.4 = 4320

11) Fabrication overhead applied

Job P = 600*12.20 = 7320

Job Q = 900*12.20 = 10980

12) calculate Unit product cost

Direct material 13000
Direct labor 21000
Molding overhead 9180
Fabrication overhead 7320
Total 50500
Unit product cost (50500/20) 2525